# ITVTON: Virtual Try-On Diffusion Transformer Based on Integrated Image and Text

### v2

@@ Abstract
{strip}[Uncaptioned image]
Figure 1: The ITVTON model is utilized to create virtual try-on images derived from the VITON-HD [4] dataset (first row) and the filtered IGPair [25] dataset (second row). For optimal visual evaluation, it is recommended to examine the images in their enlarged form.
@@ 1 Introduction
Recently, diffusion models [26, 27] have emerged as a compelling alternative for virtual try-on tasks, owing to their superior generative capabilities and enhanced flexibility. Unlike GAN-based pipelines, diffusion models [24, 23] tend to generate more stable, high-quality, and diverse outputs. They also possess strong control and optimization properties during training, making them well-suited for addressing challenges in virtual try-on. Current diffusion-based virtual try-on methods primarily adopt parallel U-shaped UNet structures [22], typically introducing a garment UNet alongside a generative UNet, with attention mechanisms bridging garment features and the person image. While this approach substantially advances virtual try-on, it also entails additional network modules and higher computational costs at both training and inference stages.
@@ 3.1 Background on Diffusion Model
The core concept of the rectified flow technique employed in FLUX is to streamline the training and inference processes by introducing a flow-matching method that enhances the efficiency of image generation. The loss function is defined as follows:
@@ 3.2 ITVTON
The ITVTON architecture, illustrated in Figure 2, aims to generate high-quality and high-fidelity virtual try-on images using a simple network structure and minimal training cost. This section outlines the components of our network architecture and the underlying architectural rationale, as well as explores efficient parameter training strategies.
@@ 3.2.1 Network Module
Text Encoder: Two text encoders (CLIP ViT-L and T5-XXL [20]) process textual prompts as input, enabling the generation of more controllable, high-quality images. Inspired by IC-LoRA [14], multi-image integrated text inputs have a positive impact on virtual try-on tasks. Therefore, we utilize multi-image integrated text (text format shown in Figure 2) as input, enhancing the controllability and quality of the generated images.
@@ 3.2.2 Architectural Inference
Specifically, consider a garment image $\mathbf{I}_{\mathbf{g}}\in\mathbb{R}^{3\times H\times W}$ and a person image $\mathbf{I}_{\mathbf{p}}\in\mathbb{R}^{3\times H\times W}$ , which are concatenated along the width dimension to generate $\mathbf{C}_{\mathbf{gp}}\in\mathbb{R}^{3\times H\times 2W}$ The corresponding binary, clothing-agnostic mask map of the person image, denoted $\mathbf{m}_{\mathbf{p}}\in\mathbb{R}^{H\times W}$, is concatenated with an all-zero mask map of the same dimensions to produce $\mathbf{m}_{\mathbf{op}}\in\mathbb{R}^{H\times 2W}$.
@@
where $\otimes$ is used for element-wise multiplication.$\mathbf{V}_{\mathbf{masked}}$ and $\mathbf{X}_{\mathbf{om}}$ undergo a packing operation to generate $\mathbf{P}_{\mathbf{masked}}\in\mathbb{R}^{64\times(\frac{H}{8\times 2}\times\frac{2W}{8\times 2})}$ and $\mathbf{P}_{\mathbf{om}}\in\mathbb{R}^{256\times(\frac{H}{8\times 2}\times\frac{2W}{8\times 2})}$, respectively, before being fed into the transformer module.The channel values are scaled by a factor of 4, while the width and height are each halved, then multiplied and compressed into a single-dimensional value.
@@ 4.3 Qualitative Comparison
We evaluated the visual quality using the VITON-HD [4] dataset. As shown in Figure 3, it demonstrates the try-on effects of various styles and patterns from the dataset, and compares the try-on variations across different methods. While other methods encounter issues such as color discrepancies and inconsistent details during virtual try-ons, ITVTON delivers more realistic and detailed outcomes owing to its simplicity and efficiency.
Additionally, we assess ITVTON on selected IGPair [25] datasets and benchmark it against other diffusion-based VTON methods. As shown in Figure 4, ITVTON demonstrates its capability to accurately recognize and model the shape of complex garments, such as a sheath dress, and seamlessly integrate them with human figures. Notably, it can generate high-quality images even in complex poses (e.g., sitting on the chair). In intricate field scenes, ITVTON effectively handles the integration of backgrounds with clothing, maintaining consistency and realism across diverse and challenging environments.
@@ 4.4 Quantitative Comparison
4.5 Ablation Studies
  1. 1.
    different trainable modules,
@@ 4.5 Ablation Studies
All Attention 0.8702 5.155 0.250 0.0557 1611.23
MM-DiT Block Attention 0.8641 5.438 0.332 0.0578 1434.93
Single-DiT Block Attention 0.8734 5.064 0.264 0.0530 1076.2
Figure 5: Train the different modules in the Transformer.
Trainable Modules: As shown in Figure 5, we evaluated three trainable modules in Transformer: 1) all attention parameters. 2) attention parameters of MM-DiT blocks. 3) attention parameters of Single-DiT blocks, with progressively fewer trainable parameters. Table 2 presents the evaluation results on the VITON-HD [4] dataset. We observe that the model trained with attention parameters within the Single-DiT block performs optimally across three metrics. The other metric, KID, shows only a slight decrease compared to the model that trains all attention parameters, while the Single-DiT model uses only 1076.2M attention parameters. Fewer training weights both reduce memory requirements and speed up training, and do not pull down performance. This demonstrates that training the attention parameters within the Single-DiT block is optimal.
@@
Multi-Image Integrated Text: Ordinary text, such as “A model is wearing a top.” and multi-image integrated text are illustrated in Figure 2. Both types of text inputs were employed during training and inference. As shown in Table 3, two metrics (SSIM and LPIPS) perform optimally when only multi-image integrated text is used during inference, although the improvement is marginal compared to the model using multi-image integrated text for both training and inference. Additionally, the other two metrics (FID and KID) are optimal for the model that utilizes multi-image integrated text for both training and inference, demonstrating that this approach is the most effective.
@@
Figure 6: Comparison of the effects of setting different guidance scales during training. The model with the guidance scale set to 30 during inference and the guidance scale set to 2 or 30 during training does not perform well when generating garments with text or complex patterns. In contrast, the model with the guidance scale set to 3.5 during training performs well.
Figure 7: Comparison of the effects of setting different guidance scales during inference. Setting the guidance scale to 3.5 for training and 30 or 45 for inference yields the best performance. If set to other values, issues such as color discrepancies and pattern disruption may occur.
Guidance Scale: The guidance scale is crucial in both the training and inference processes of the FLUX model. First, adjust the guidance scale setting to 30 for inference, as shown in Table 4. The metrics FID, KID, and LPIPS achieve optimal results with a guidance scale setting of 3.5 during training, while the SSIM metric is slightly lower compared to the model with a guidance scale of 30 during training. This suggests that setting the guidance scale to 3.5 during training yields better performance. As shown in Figure 6, models with the guidance scale set to 2 or 30 during training do not perform well when generating garments with repeated or intricate patterns, while models with the guidance scale set to 3.5 can accurately reproduce repeated patterns or images. Next, set the guidance scale to 3.5 for training and evaluate the effects of different guidance scale settings during inference. As shown in Figure 7, setting the guidance scale to 30 or 45 during inference results in the best restoration of color and pattern. For example, setting the guidance scale to 0 during inference results in color discrepancies and broken patterns. In summary, setting the guidance scale to 3.5 for training and 30 for inference yields the best inference results.
4.6 Constraints and Social Impacts
Although our model yields high-quality try-on results, it has certain limitations. The accuracy of the try-on process is highly dependent on the precision of the input mask; inaccurate masking can significantly degrade the quality of the generated output. For instance, if the mask fails to fully cover the person’s garment, the try-on results will be adversely affected by exposed areas of the garment. Additionally, character features such as hairstyles, jewelry, muscle contours, and tattoos within the masked area may undergo alteration during the generation process. These limitations underscore the necessity for precise preprocessing and further refinement of the model to ensure the consistency and reliability of the output.
From a societal perspective, ITVTON can help users to effectively visualize themselves in a particular outfit, and therefore our approach must be used responsibly, requiring users to be careful to protect image ownership from malicious manipulation, such as unethically modifying personal images.
@@ 5 Conclusion
In this paper, we introduced ITVTON, a simple yet effective virtual try-on diffusion transformer model with 1,076.2M trainable parameters. ITVTON spliced garment and person images along the width dimension, taken integrated image-text descriptions as input, and required only fine-tuning to generate high-quality, high-fidelity virtual try-on images. Extensive experiments have demonstrated that ITVTON could achieve superior qualitative and quantitative results compared to state-of-the-art methods, while maintaining a simple and efficient network architecture. These findings also suggested that ITVTON would hold significant potential for widespread application in virtual try-on technology.

### v3

@@ Abstract
Figure 1: The ITVTON model is utilized to create virtual try-on images derived from the VITON-HD [4] dataset (row 1) and the filtered IGPair [23] dataset (row 2). For optimal visual evaluation, it is recommended to examine the images in enlarged form.
@@ 1 Introduction
Recently, diffusion models [25, 24] have emerged as a compelling alternative for virtual try-on tasks, owing to their superior generative capabilities and enhanced flexibility. Unlike GAN-based pipelines, diffusion models [26] tend to generate more stable, high-quality, and diverse outputs. They also possess strong control and optimization properties during training, making them well-suited for addressing challenges in virtual try-on. Current diffusion-based virtual try-on methods primarily adopt parallel U-shaped UNet structures [22], typically introducing a garment UNet alongside a generative UNet, with attention mechanisms bridging garment features and the person image. While this approach substantially advances virtual try-on, it also entails additional network modules and higher computational costs at both training and inference stages.
@@ 3.1 Background on Diffusion Model
The core concept of the rectified flow technique employed in FLUX is to streamline the training and inference processes by introducing a flow-matching method that enhances the efficiency of image generation. The loss function is defined as
@@ 3.2 ITVTON
The ITVTON architecture, illustrated in Fig. 2, aims to generate high-quality and high-fidelity virtual try-on images using a simple network structure and minimal training cost. This section outlines the components of our network architecture and the underlying architectural rationale, as well as explores efficient parameter training strategies.
@@ 3.2.1 Network Module
Text Encoder: Two text encoders (CLIP ViT-L and T5-XXL [20]) process textual prompts as input, enabling the generation of more controllable, high-quality images. Inspired by IC-LoRA [14], multi-image integrated text inputs have a positive impact on virtual try-on tasks. Therefore, we utilize multi-image integrated text (text format shown in Fig. 2) as input, enhancing the controllability and quality of the generated images.
@@ 3.2.2 Architectural Inference
Specifically, consider a garment image $\mathbf{I}_{\mathbf{g}}\in\mathbb{R}^{3\times H\times W}$ and a person image $\mathbf{I}_{\mathbf{p}}\in\mathbb{R}^{3\times H\times W}$ , which are concatenated along the width dimension to generate $\mathbf{C}_{\mathbf{gp}}\in\mathbb{R}^{3\times H\times 2W}$ . The corresponding binary, clothing-agnostic mask map of the person image, denoted $\mathbf{m}_{\mathbf{p}}\in\mathbb{R}^{H\times W}$, is concatenated with an all-zero mask map of the same dimensions to produce $\mathbf{m}_{\mathbf{op}}\in\mathbb{R}^{H\times 2W}$.
@@
[…]
@@ 4.3 Qualitative Comparison
We evaluated the visual quality using the VITON-HD [4] dataset. As shown in Fig. 3, it demonstrates the try-on effects of various styles and patterns from the dataset, and compares the try-on variations across different methods. While other methods encounter issues such as color discrepancies and inconsistent details during virtual try-ons, ITVTON delivers more realistic and detailed outcomes owing to its simplicity and efficiency.
Additionally, we assess ITVTON on selected IGPair [23] datasets and benchmark it against other diffusion-based VTON methods. As shown in Fig. 5, ITVTON demonstrates its capability to accurately recognize and model the shape of complex garments, such as a sheath dress, and seamlessly integrate them with human figures. Notably, it can generate high-quality images even in complex poses (e.g., sitting on the chair). In intricate field scenes, ITVTON effectively handles the integration of backgrounds with clothing, maintaining consistency and realism across diverse and challenging environments.
@@ 4.5 Ablation Studies
    All Attention 0.8702 5.155 0.250 0.0557 1611.23
    MM-DiT Block Attention 0.8641 5.438 0.332 0.0578 1434.93
    Single-DiT Block Attention 0.8734 5.064 0.264 0.0530 1076.2
@@
Trainable Modules: We evaluated three trainable modules in Transformer: 1) all attention parameters. 2) attention parameters of MM-DiT blocks. 3) attention parameters of Single-DiT blocks, with progressively fewer trainable parameters. Table 2 presents the evaluation results on the VITON-HD [4] dataset. We observe that the model trained with attention parameters within the Single-DiT block performs optimally across three metrics. The other metric, KID, shows only a slight decrease compared to the model that trains all attention parameters, while the Single-DiT model uses only 1076.2M attention parameters. Fewer training weights both reduce memory requirements and speed up training, and do not pull down performance. This demonstrates that training the attention parameters within the Single-DiT block is optimal.
Multi-Image Integrated Text: Ordinary text, such as “A model is wearing a top.”. Multi-image integrated text is shown in Fig. 2. Both types of text inputs were employed during training and inference. As shown in Table 3, two metrics (SSIM and LPIPS) perform optimally when only multi-image integrated text is used during inference, although the improvement is marginal compared to the model using multi-image integrated text for both training and inference. Additionally, the other two metrics (FID and KID) are optimal for the model that utilizes multi-image integrated text for both training and inference, demonstrating that this approach is the most effective.
Figure 5: Comparison of the effects of setting different guidance scales during training. The model with the guidance scale set to 30 during inference and the guidance scale set to 2 or 30 during training does not perform well when generating garments with text or complex patterns.
Figure 6: Comparison of the effects of setting different guidance scales during inference.
Guidance Scale: The guidance scale is crucial in both the training and inference processes of the FLUX model. First, adjust the guidance scale setting to 30 for inference, as shown in Table 4. The metrics FID, KID, and LPIPS achieve optimal results with a guidance scale setting of 3.5 during training, while the SSIM metric is slightly lower compared to the model with a guidance scale of 30 during training. This suggests that setting the guidance scale to 3.5 during training yields better performance. As shown in Fig. 5, models with the guidance scale set to 2 or 30 during training do not perform well when generating garments with repeated or intricate patterns, while models with the guidance scale set to 3.5 can accurately reproduce repeated patterns or images. Next, set the guidance scale to 3.5 for training and evaluate the effects of different guidance scale settings during inference. As shown in Fig. 6, setting the guidance scale to 30 or 45 during inference results in the best restoration of color and pattern. For example, setting the guidance scale to 0 during inference results in color discrepancies and broken patterns. In summary, setting the guidance scale to 3.5 for training and 30 for inference yields the best inference results.
@@ 5 Conclusion
In this paper, we introduced ITVTON, a simple yet effective virtual try-on diffusion transformer model with 1,076.2M trainable parameters. ITVTON spliced garment and person images along the width dimension, taken integrated image-text descriptions as input, and required only fine-tuning to generate high-quality, high-fidelity virtual try-on images.
Extensive experiments have demonstrated that ITVTON could achieve superior qualitative and quantitative results compared to state-of-the-art methods, while maintaining a simple and efficient network architecture. These findings also suggested that ITVTON would hold significant potential for widespread application in virtual try-on technology.